# Reproducible Machine Learning-based Voice Pathology Detection: Introducing the Pitch Difference Feature

### v2

@@ 4.1 Pitch Difference
4.2 NaN Feature
For eight recordings, specifically 492-a_n.wav, 719-a_n.wav, 720-a_n.wav, 915-a_n.wav, 1338-a_n.wav, 1407-a_n.wav, 1716-a_n.wav, and 2235-a_n.wav, the algorithm fails to extract information about f0 and therefore, the feature values for $\mu_{f_{0}}$, $\Delta f_{0}$, $\sigma_{f_{0}}$, jitta, and shim are NaN. This is probably caused by the dominance of the disharmonic part for serious cases of voice pathology. We replace the NaN values with zero and add a new binary feature that takes the value of 1 for the occurrence of NaN values. As this inability to detect the fundamental frequency is an important information about the recording, we consider introducing this feature a legitimate approach. Note that these eight recordings represent 2 female and 6 male subjects in our dataset, and all of them suffer from some form of voice pathology according to the database records.
4.3 Feature Sets
In the extracted features, there are significant differences in some feature values between male and female patients. This is likely caused by the higher pitch of the female voice compared to the male voice, altering characteristics such as f0 (see Figure 6). Since several features are dependent on f0, even if indirectly, and the patient’s biological sex is known during examination, we treat the data as two separate datasets and train two separate classifiers, one for each sex. To test this approach, we also compute the results for both sexes together.
@@ 4.3 Feature Sets
5 Data Augmentation
In order to address the issue of class imbalance in the utilized dataset, which leads to classifiers with high recall and low specificity, we employ the k-means SMOTE algorithm [43]. This technique is specifically applied to the training set to enhance the model’s ability to learn from minority class instances and thus improve its predictive performance.
@@ 5 Data Augmentation
6 Machine Learning Models
We perform binary classification using NB, KNN, DT, random forest (RF), AdaBoost, and SVM. Given the relatively high dimensionalilty of the feature sets with a limited number of samples, deep learning models could be challenging to apply effectively. Therefore, we believe that using traditional ML models is a suitable approach and aligns with the findings from the existing research.
All evaluated ML models were implemented using the scikit-learn v1.5.2 library [46]. The list of tuned hyperparameters and their corresponding values for each classifier is presented in Table 3. The naming of hyperparameters in the tables in the following text corresponds to the naming of function parameters in scikit-learn.
Table 3: ML classifier hyperparameters tested for optimization
Classifier Hyperparameter Tested values
SVM kernel "rbf", "poly"
degree ("poly" only) 2, 3, 4, 5, 6
gamma 0.5, 0.1, 0.05, 0.01, 0.005, 0.001, "auto"
@@ 6 Machine Learning Models
weights "uniform", "distance"
DT criterion "gini", "entropy", "log_loss"
splitter "best", "random"
@@
max_features "sqrt", "log2"
RF criterion "gini"
@@
7 Validation of Results
It is crucial to establish robust performance metrics for our model to accurately assess its capability in distinguishing between healthy and pathological patients. Equally important is to ensure that the reported validation results are robust, minimizing the influence of randomness to guarantee the reliability and consistency of our conclusions, as well as the possibility to independently reproduce our findings.
7.1 Used Metrics
The imbalance between healthy and pathological samples in the dataset (407 healthy females, 541 females with pathologies, 252 healthy males and 436 males with pathologies) introduces bias into commonly used metrics such as accuracy, F1 score, precision, and negative predictive value [47]. Especially accuracy can dangerously show overoptimistic results and provide misleading information [48].
To address this issue and to provide metrics that reflect class imbalance, we evaluate the performance of the ML algorithms using sensitivity (Equation 2), specificity (Equation 3), unweighted average recall (Equation 5), geometric mean (Equation 4), and bookmaker informedness (Equation 6). As all these metrics provide information about the successful classification, we also evaluate MCC (Equation 7), which, although not entirely unbiased, has a smaller bias compared to accuracy and also takes misclassification into account [47].
@@ 7.1 Used Metrics
Note that bookmakers informedness (BM), UAR, and geometric mean (GM) give the same weight to sensitivity and specificity, regardless of the distribution of positive and negative samples in the dataset. BM and GM penalize the performance on the minority class more copmared to UAR, resulting in 0 values for 0 sensitivity (or specificity).
7.2 Validation Approach
Due to the limited size of the dataset and its imbalance, we employ stratified 10-fold cross-validation during the grid search, which should lead to less biased results of classifier performances [49]. Only the training folds are augmented using the k-means SMOTE-based algorithm (see Section 5) in each iteration. Then, each feature in the training set is scaled to an interval $[0,1]$ with the min-max scaler. The parameters obtained for the scaling of the training folds are used to scale the validation fold to avoid potential data leakage.
@@ 7.2 Validation Approach
1:for each classifier type do
2:    load results of all classifiers
@@
5:    for for each classifier with corresponding feature set and hyperparameters from previous step do
@@
20:    select the classifier with corresponding feature set and hyperparameters that has the highest average MCC as the best performing classifier
@@
This process yields ten values for each performance metric. We then calculate the mean value of each metric across the folds, following standard cross-validation practice.
@@
After completing the grid search, we identify the top 1000 performing classifier configurations for each classifier type based on MCC. To account for the variance in MCC introduced by the cross-validation splits, we perform 100 times repeated stratified 10-fold cross-validation for these configurations. This allows us to estimate the average metrics for each classifier model and their corresponding standard deviations more reliably. The validation of the best results is shown in Algorithm 1.
First, we aggregate the results to get the average performance as well as the standard deviation for each feature set and classifier combination (averaging out different configurations for the classifier) to find the best performing feature set for each type of classifier. Next, we sort the results according to MCC and select the best performing feature set for each classifier based on the maximum of MCC.
7.3 Ensuring Reproducibility
To maintain the reproducibility of our findings, we implement multiple mechanisms so that all computations provide the same results and everyone can verify that our results were produced by the code we supply. The main problem arises from unspecified license of the SVD. Therefore, we cannot share the feature sets, and we implement methods based on SHA256 checksums to validate input data and intermediate results. To make the use of methodology easier, we include a file named svd_information.csv in the code repository, which contains metadata for each recording session, such as age, pathology etc. The full list of files downloaded for this work is included in the repository along with SHA256 checksums.
@@ 8 Results & Discussion
As we mentioned and explained in Section 7, we do not present the accuracy score as the data used for training and evaluation of the dataset is moderately imbalanced. Instead, we provide an alternative in the form of MCC, UAR, GM, and BM. However, if needed, accuracy can still be found in raw results reported in the repository.
8.1 Top Performing Classifiers
The results for the best performing classifiers for females, males and both sexes including classifier type are presented in Table 4. The configurations of the feature sets reaching the best results for the classifier are in Table 5 for females, in Table 6 for males, and in Table 7 for both sexes. Note, that a zero standard deviation for NB corresponds to a subtle effect of Lidstone smoothing, which is manifested only in higher decimal places. Moreover, the results are obtained by 100 repetitions of stratified 10-fold cross-validation.
Table 4: Best results reached for each classifier type
Sex Classifier MCC SEN SPE GM UAR BM
@@ 8.1 Top Performing Classifiers
Table 5: Configuration of the best model and feature set for each classifier type - women
Classifier AdaBoost DT NB KNN RF SVM
@@
Table 6: Configuration of the best model and feature set for each classifier type - men
Classifier AdaBoost DT NB KNN RF SVM
@@
Table 7: Configuration of the best model and feature set for each classifier type - both
Classifier AdaBoost DT NB KNN RF SVM
@@
Based on the MCC score, the best performing male classifier is SVM. For the female dataset, the best performing model is AdaBoost, while the second best is RF. It is worth mentioning that the SVM model performed in classification of female dataset just slightly worse than both AdaBoost and RF model. For both sexes, the best performing model is SVM. The female classifiers perform better in general, which might be partially affected by the higher sample size and better ratio between the pathological and healthy samples. The top four female classifiers outperformed the top male and both sexes classifiers in MCC.
Our proposed feature, the pitch difference, appears in the top performing classifiers, being more prominent in the female feature sets. Specifically, it is utilized by the best performing female AdaBoost, RF, SVM and KNN classifiers and by the best performing classifier SVM for both sexes. For male classifiers, the pitch difference was used only in RF classifier, which was the third best classifier. The NaN feature was used by the best female classifiers only, namely the AdaBoost and RF classifiers.
Feature sets with 13 MFCC features were not present among the best performing classifiers and 20 MFCC features were used only by the female and male AdaBoost classifiers. The variances of MFCC were used by the male and female AdaBoost classifier. Notably, the top female AdaBoost classifier utilized almost complete feature set, where only formants, spectral flatness and zero-crossing rate were omitted. Additionally, the entropy and LFCC were used by the top female, male and both-sex classifiers. Generally, all features, except 13 MFCC were used at least by one type of the best performing classifier.
@@ 8.2 Top Performing Feature Sets
Table 8: Best average feature set performance for each classifier - male patients
Classifier SVM KNN NB DT RF AdaBoost
@@
Table 9: Best average feature set performance for each classifier - female patients
Classifier SVM KNN NB DT RF AdaBoost
@@
Table 10: Best average feature set performance for each classifier - both sexes
Classifier SVM KNN NB DT RF AdaBoost
@@
The best performing feature sets for each classifier type are selected by averaging their MCC across all hyperparameter configurations (see Section 6). The best performing feature sets in combination with each classifier type are in Table 9 for females, in Tables 8 for males, and in Table 10 for both sexes.
It is important to note that the comparison of feature sets results across all hyperparameter configurations was conducted without repeated cross-validation, therefore, the performance should not be compared between different classifiers and to the best performing feature sets as the results might be influenced by the split to individual folds.
The variances of MFCC were absent completely. LFCC were not used in female feature sets and spectral roll-off were not used in male feature sets. The remaining features, including pitch difference and NaN feature were present at least once.
8.3 General Discussion
To support our preference for MCC instead of accuracy, we estimate the biases for each metric for the best classifier in Table 12, according to [47].
Table 11: Related works on full SVD
Article UAR Notes
[4] 73.3% [^1] Value computed from other metrics reported in the article.
[5] 77.86% [^1] Value computed from other metrics reported in the article.
[6] 70.68% [^1] Value computed from other metrics reported in the article.
[7] Inconsistent results, UAR unknown
[8] 72.32% [^1] Value computed from other metrics reported in the article.
[9] 84.97%
[10] 88.75% [^1] Value computed from other metrics reported in the article. Potential data-leakage.
[12] 84.55%
Our approach 85.22% [^2] Best results for each sex weighted by number of samples of each sex
Table 12: Estimated mean bias of each metric for best classifier for each sex
@@ 8.3 General Discussion
MCC (normalized)  [^3] Normalized MCC is transformation of MCC to the interval [0,1] 0.8388 -0.0170 0.8378 -0.0091 0.8463 -0.0029
The majority of the reported works excluded data based on individual pathologies. While this may improve the classifier’s performance and even allow multimodal classification of individual pathologies, we strongly believe this approach actually limits the applicability. In clinical practice, these classifiers could not be utilized unless the excluded pathologies were also excluded from the possible diagnoses for the examined patients, which may prove impractical for clinical use.
None of the investigated studies reported handling the potential data leakage stemming from training classifiers on datasets containing multiple recordings from the same patients. As it is a common knowledge that ML models can learn patterns of individual patients, given enough input, we assume many results of the works in ML-based voice pathology detection might be overestimated due to this error. Therefore, we applied a method to exclude duplicate data based on patient identity. Our proposed approach does not discriminate against either healthy or pathological data and does not lead to overoptimistic results.
The results we managed to reach with our methodology are on par with or surpass the reported performance in Section 2 and are summarized in Table 11. Our result is the combination of the best sex-aware classifiers with results weighted by the number of subject of each sex. To our best knowledge, this is the first paper on voice pathology detection combining SVD and ML methods, that is fully reproducible and conforms to the REFORMS practices [52]. We provide the filled in REFORMS checklist in Appendix A.
All features, except the three indicated in this paragraph, are obtained from voice recordings and are widely used in the models for voice pathology detection (see subsection Section 2). We consider age legitimate as a feature, as other acoustic features depend on the age of the speaker, for example, changes in speaking fundamental frequency with aging [53]. The pitch difference (see subsection Section 4.1) is our proposed feature indicating the change of fundamental frequency during the recording and is extracted in a similar way as other considered features. The NaN feature reflects the fact that it was not possible to estimate the fundamental frequency for the patient (as described in Subsection 4.3). As the fundamental frequency is considered one of the dominant features in voice pathology diagnosis, we regard NaN feature as a legitimate approach.
@@ 8.4 Limitations
We are aware of several limitations our work is subject to. First, our model was tested using SVD only. The used database does not fully reflect the general population, especially in the proportion of healthy and pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as the only viable source of data, we cannot extrapolate its performance outside of this dataset. As the data was recorded in a controlled environment, we can assume our classifiers might not be able to perform as well with datasets that are recorded during different conditions.
Moreover, we limit our research only to individuals who are 18 years old and older. Another noteworthy limitation was the available computational capacity, which led to careful decision of the classification algorithms, hyperparameter space and selected features we drew on throughout our work.

### v3

@@ 3.2.1 Pitch Difference
3.2.2 NaN Feature
For eight recordings, specifically 492-a_n.wav, 719-a_n.wav, 720-a_n. wav, 915-a_n.wav, 1338-a_n.wav, 1407-a_n.wav, 1716-a_n.wav, and 2235-a_n.wav, the algorithm extracting the f0, implemented in parselmouth library, fails and therefore, the feature values for $\mu_{f_{0}}$, $\Delta f_{0}$, $\sigma_{f_{0}}$, jitta, and shim are NaN. This is probably caused by the dominance of the disharmonic part for serious cases of voice pathology. We replace the NaN values with zero and add a new binary feature that takes the value of 1 for the occurrence of NaN values. As this inability to detect the fundamental frequency is an important information about the recording, we consider introducing this feature a legitimate approach. Note that these eight recordings represent 2 female and 6 male subjects in our dataset, and all of them suffer from some form of voice pathology according to the database records.
3.2.3 Feature Sets
In the extracted features, there are significant differences in some feature values between male and female patients. This is likely caused by the higher pitch of the female voice compared to the male voice, altering characteristics such as f0 (see Figure 6). Since several features are dependent on f0, even if indirectly, and the patient’s biological sex is known during examination, we treat the data as two separate datasets and train two separate classification models, one for each sex. To verify this approach, we also train models for both sexes together.
@@ 3.2.3 Feature Sets
3.3 Data Augmentation
In order to address the issue of class imbalance in the utilized dataset, which leads to models with high recall and low specificity, we employ the k-means SMOTE algorithm [51]. This technique is specifically applied to the training set to enhance the model’s ability to learn from minority class instances and thus improve its predictive performance.
@@ 3.3 Data Augmentation
3.4 Machine Learning Algorithms
We perform binary classification using NB, KNN, DT, random forest (RF), AdaBoost, and SVM. Given the relatively high dimensionality of the feature sets with a limited number of samples, deep learning models could be challenging to apply effectively. Therefore, we believe that using traditional ML algorithms is a suitable approach and aligns with the findings from the existing research.
All evaluated ML algorithms were implemented using the scikit-learn v1.5.2 library [54]. The list of hyperparameters tuned using grid search along with their corresponding values for each algorithm is presented in Table 3. The naming of hyperparameters in the tables in the following text corresponds to the naming of function parameters in scikit-learn.
Table 3: ML algorithm hyperparameters for grid search
Algorithm Hyperparameter Tested values
SVM kernel ”rbf”, ”poly”
degree (”poly” only) 2, 3, 4, 5, 6
gamma 0.5, 0.1, 0.05, 0.01, 0.005, 0.001, ”auto”
@@ 3.4 Machine Learning Algorithms
weights ”uniform”, ”distance”
DT criterion ”gini”, ”entropy”, ”log_loss”
splitter ”best”, ”random”
@@
max_features ”sqrt”, ”log2”
RF criterion ”gini”
@@
3.5 Validation of Results
It is crucial to establish robust performance metrics to accurately assess model capability in distinguishing between healthy and pathological patients. Equally important is to ensure that the reported validation results are robust, minimizing the influence of randomness to guarantee the reliability and consistency of our conclusions, as well as the possibility to independently reproduce our findings.
3.5.1 Used Metrics
The imbalance between healthy and pathological samples in the dataset (407 healthy females, 541 females with pathologies, 252 healthy males and 436 males with pathologies) introduces bias into commonly used metrics such as accuracy, F1 score, precision, and negative predictive value [55]. Especially, accuracy can dangerously show overoptimistic results and provide misleading information [56].
To address this issue and to provide metrics that reflect class imbalance, we evaluate the model performance using sensitivity (Equation 2), specificity (Equation 3), UAR (Equation 5), geometric mean (GM) (Equation 4), and bookmakers informedness (BM) (Equation 6). As all these metrics provide information about the successful classification, we also evaluate MCC (Equation 7), which, although not entirely unbiased, has a smaller bias compared to accuracy and also takes misclassification into account [55].
@@ 3.5.1 Used Metrics
Note that BM, UAR, and GM give the same weight to sensitivity and specificity, regardless of the distribution of positive and negative samples in the dataset. BM and GM penalize the performance on the minority class more compared to UAR, resulting in 0 values for 0 sensitivity (or specificity).
3.5.2 Validation Approach
Due to the limited size of the dataset and its imbalance, we employ stratified 10-fold cross-validation during the grid search, which should lead to less biased results of model performances [57]. Only the training folds are augmented using the k-means SMOTE-based algorithm (see Section 3.3) in each iteration. Then, each feature in the training set is scaled to an interval $[0,1]$ with the min-max scaler. The parameters obtained for the scaling of the training folds are used to scale the validation fold to avoid potential data leakage.
@@ 3.5.2 Validation Approach
1:for each ML algorithm do
2:    load results of all models
@@
5:    for each model with corresponding feature set from previous step do
@@
20:    select the model with corresponding feature set that has the highest average MCC as the best performing model
@@
This process yields ten values for each performance metric. We then calculate the mean value of each metric across the folds, following standard cross-validation practices.
@@
After completing the grid search, we identify the top 1000 performing models for each ML algorithm based on MCC. To account for the variance in MCC introduced by the cross-validation splits, we perform 100 times repeated stratified 10-fold cross-validation for these models. This allows us to estimate the average metrics for each classification model and their corresponding standard deviations more reliably. The validation of the best results is shown in Algorithm 1.
Moreover, we determine the top performing feature sets for each ML algorithm. First, we aggregate the results from the grid search to get the average performance as well as the standard deviation for each feature set and ML algorithm combination. Then, we select the best feature set for each ML algorithm based on the highest value of MCC.
3.5.3 Ensuring Reproducibility
To maintain the reproducibility of our findings, we implement multiple mechanisms so that all computations provide the same results and everyone can verify that our results were produced by the code we supply. The main problem arises from the unspecified license of the SVD. Therefore, we cannot share the feature sets, and we implement methods based on SHA256 checksums to validate input data and intermediate results. To make the use of methodology easier, we include a file named svd_information.csv in the code repository, which contains metadata for each recording session, such as age, pathology, etc. The full list of files downloaded for this work is included in the repository along with SHA256 checksums.
@@ 4 Results
As we mentioned and explained in Section 3.5, we do not present the accuracy score as the data used for training and evaluation of the dataset is moderately imbalanced. Instead, we provide an alternative in the form of MCC, UAR, GM, and BM. However, if needed, accuracy can still be found in raw results reported in the repository.
4.1 Top Performing Models
The results for the best performing models (see step 20 of Algorithm 1) for females, males and both sexes are presented in Table 4. The corresponding features and hyperparameter setting are in Table 5 for females, in Table 6 for males, and in Table 7 for both sexes. Note, that a zero standard deviation for NB corresponds to a subtle effect of Lidstone smoothing, which is manifested only in higher decimal places.
Table 4: Top model performance for each ML algorithm
Sex Algorithm MCC SEN SPE GM UAR BM
@@ 4.1 Top Performing Models
Table 5: Top model – feature set configuration for each ML algorithm — females
Algorithm AdaBoost DT NB KNN RF SVM
@@
Table 6: Top model – feature set configuration for each ML algorithm — males
Algorithm AdaBoost DT NB KNN RF SVM
@@
Table 7: Top model – feature set configuration for each ML algorithm — both
Algorithm AdaBoost DT NB KNN RF SVM
@@
Based on the MCC score, the best model for males is SVM. For the female dataset, the best performing model is AdaBoost, while the second best is RF. It is worth mentioning that the SVM model performed in classification of female dataset just slightly worse than both AdaBoost and RF model. For both sexes, the best performing model is SVM. The models for females perform better in general, which might be partially affected by the higher sample size and better ratio between the pathological and healthy samples. The top four female models outperformed the top male models and models including both sexes in MCC.
Our proposed feature, the pitch difference, appears in the feature sets of the best performing models, being more prominent in the female feature sets. Specifically, it is utilized by the best performing female AdaBoost, RF, SVM, and KNN models and by the best performing SVM model for both sexes. Regarding classification of males, the pitch difference was used only in RF model, which was the third best model. The NaN feature was used by the best female models only, namely the AdaBoost and RF models.
Feature sets with 13 MFCC features were not present among the best performing models and 20 MFCC features were used only by the female and male AdaBoost models. The variances of MFCC were used by the male and female AdaBoost models. Notably, the top female AdaBoost models utilized almost complete feature set, where only formants, spectral flatness, and zero-crossing rate were omitted. Additionally, the entropy and LFCC were used by the top female, male, and both-sex models. Generally, all features, except 13 MFCC, were used at least by one top performing model.
@@ 4.2 Top Performing Feature Sets
Table 8: Top performing feature set for each ML algorithm — males
Algorithm SVM KNN NB DT RF AdaBoost
@@
Table 9: Top performing feature set for each ML algorithm — females
Algorithm SVM KNN NB DT RF AdaBoost
@@
Table 10: Top performing feature set for each ML algorithm — both
Algorithm SVM KNN NB DT RF AdaBoost
@@
The best performing feature sets for each ML algorithm are selected by averaging their MCC across all hyperparameter configurations (see Section 3.4). The best performing feature sets in combination with each ML algorithm type are in Table 9 for females, in Tables 8 for males, and in Table 10 for both sexes.
Note that for each ML algorithm, a different number of models was trained due to the different number of possible hyperparameter combinations. Moreover, the training was conducted on a single 10-fold cross-validation split which has a different influence on the final performance of each ML algorithm. Therefore, the performance between models is not comparable.
The results show that the variances of MFCC were absent completely in the best performing feature sets. LFCC were not used in female feature sets and spectral roll-off was not used in male feature sets. The remaining features, including pitch difference and NaN feature, were present at least once.
5 Discussion
Table 11: Estimated mean bias of each metric for the best model for each sex
@@ 5 Discussion
MCC (normalized) 3 0.8388 -0.0170 0.8378 -0.0091 0.8463 -0.0029
  • 3
    Normalized MCC is transformation of MCC to the interval [0,1]
To support our preference for MCC instead of accuracy, we estimate the biases for each metric for the best model in Table 11, according to [55].
Many studies reported in Related Works section exclude data based on individual pathologies. While this may improve model performance and even allow multimodal classification of individual pathologies, we strongly believe this approach actually limits the applicability. In clinical practice, these models could not be utilized unless the excluded pathologies were also excluded from the possible diagnoses for the examined patients, which may prove impractical for clinical use.
None of the investigated studies reported handling the potential data leakage stemming from training models on datasets containing multiple recordings from the same patients. As there is a possibility that models can learn patterns of individual patients, given enough input, we assume many results of the works in ML-based voice pathology detection might be overestimated due to this error. Therefore, we applied a method to exclude duplicate data based on patient identity. Our proposed approach does not discriminate against either healthy or pathological data and does not lead to overoptimistic results.
Our result is the combination of the best sex-aware models with results weighted by the number of subject of each sex. To our best knowledge, this is the first paper on voice pathology detection combining SVD and ML methods, that is fully reproducible and conforms to the REFORMS practices [60]. We provide the filled in REFORMS checklist in A.
All features, except the three indicated in this paragraph, are obtained from voice recordings and are widely used in the models for voice pathology detection (see subsection Section 2). We consider age legitimate as a feature, as other acoustic features depend on the age of the speaker — for example, changes in speaking fundamental frequency with aging [61]. The pitch difference (see subsection Section 3.2.1) is our proposed feature indicating the change of fundamental frequency during the recording and is extracted in a similar way as other considered features. The NaN feature reflects the fact that it was not possible to estimate the fundamental frequency for the patient (as described in Subsection 3.2.3). As the fundamental frequency is considered one of the dominant features in voice pathology diagnosis, we regard NaN feature as a legitimate approach.
@@ 5.1 Limitations
We are aware of several limitations our work is subject to. First, our models were tested using SVD only. The used database does not fully reflect the general population, especially in the proportion of healthy and pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as the only viable source of data, we cannot extrapolate its performance outside of this dataset. As the data was recorded in a controlled environment, we can assume our models might not be able to perform as well with datasets that are recorded under different conditions.
Moreover, we limit our research only to individuals who are 18 years old and older. Another noteworthy limitation was the available computational capacity, which led to careful decision of the ML algorithms, hyperparameter space, and selected features we drew on throughout our work.
We also acknowledge that many voice and speech disorders extend beyond phonatory function, affecting articulation, prosody, and temporal dynamics, which are best analyzed through continuous speech rather than isolated vowel production. Neuromotoric diseases often present with abnormalities in articulation rate, syllable duration, coarticulation patterns, and prosodic modulation. These features cannot be fully evaluated through sustained vowel phonation alone, as they require an analysis of connected speech, where natural variations in stress, intonation, and fluency become more apparent [62, 63] in the context of the four-level model and its implications for understanding the pathophysiology underlying apraxia of speech and other motor speech disorders [64].
In addition, deficits in speech motor control, imprecise consonant production, and irregularities in rhythm, which are key characteristics of certain neurological disorders, are best observed in spontaneous or structured speech tasks rather than in isolated vowel production. Including continuous speech in voice pathology detection frameworks enables a more comprehensive assessment of speech intelligibility, temporal variability, and segmental articulation, ultimately enhancing diagnostic accuracy and the clinical relevance of machine learning models. Acknowledging these limitations, future research should consider incorporating both sustained phonation and continuous speech samples into pathology detection systems to improve sensitivity to a wider range of voice and speech disorders.